OpenStudy (anonymous):

find the exact value of cos(-75) and sin(-75) justify your answer

OpenStudy (mertsj):

\[\cos (-30+-45)= \cos (-30)\cos (-45)-\sin (-30)\sin (-45)=\frac{\sqrt{3}}{2}\frac{\sqrt{2}}{2}-\frac{-1}{2}\frac{-\sqrt{2}}{2}=\]

OpenStudy (mertsj):

\[\frac{\sqrt{6}}{4}-\frac{\sqrt{2}}{4}\]

OpenStudy (mertsj):

\[\sin (-30+-45)=\sin (-30)\cos (-45)+\cos (-30)\sin (-45)=\frac{-1}{2}\frac{\sqrt{2}}{2}+\frac{\sqrt{3}}{2}(\frac{-\sqrt{2}}{2})=\]

OpenStudy (mertsj):

\[\frac{-\sqrt{2}}{4}-\frac{\sqrt{6}}{4}\]

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

yw